An investor, sold 400 shares short of XYZl, Inc. at $42 a share. The price of the stock subsequently fell to $38 before rising to $50 at which time investor covered the position (that is, closed the short position). What was the percentage gain or loss on this investment?
Answer #1

Given that,

An investor, sold 400 shares short of XYZl, Inc. at $42 a share

Than he covered the position at when stock fall to $38

So, total profit from short sell is

Profit = (selling price - buying price) = (42 - 38) = $4

So, percentage gain = profit/selling price = 4/42 = 9.52%